What is the molarity of 87% phosphoric acid solution? (Density of the solution is 1.46 g/mL). Please provide a detailed explanation.
Answer #1

87% Phosphoric acid means 87 g. of H3PO4 in 100 g of phosphoric acid solution.

Thus from density, Volume=Mass/Density = 87/1.46 = 59.58 ml

No. of moles = Mass/Molar Mass

Molar Mass Of H3PO4 is 98 g.

No. of moles = 87/98 = 0.88 moles

Molarity =(0.88/59.58) * 1000 = 14.77 M